OpenStudy (anonymous):

I DON'T UNDERSTAND PLEASE HELP ASAP!!!! Can someone please 1st explain the question to see if I understand what I am supposed to do and then show step by step I would really appreciate it! An equation is shown below: 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (cggurumanjunath):

put a value for x and you get corresponding value for y

OpenStudy (cggurumanjunath):

9x = -2 + y or y = 9x + 2

OpenStudy (cggurumanjunath):

y = 9x + 2

OpenStudy (cggurumanjunath):

put x = 1 y = 9 (1) + 2 = 9 + 2 = 11

OpenStudy (cggurumanjunath):

put x = 2 y = 9(2)+2 =18+2=20

OpenStudy (anonymous):

ok hold on I will be right back going to get me some paper and pen so I can take notes thanks for your help

OpenStudy (cggurumanjunath):

put x = 3 y = 9(3) +2= 27 +2 =29

OpenStudy (anonymous):

I'm back writing this now

OpenStudy (cggurumanjunath):

(1,11) , (2,20) and (3,29) are three different solutions !

OpenStudy (cggurumanjunath):

part A : y = 9x + 2

OpenStudy (cggurumanjunath):

part B : (1,11) , (2,20) and (3,29) are three different solutions !

OpenStudy (anonymous):

Ok I am right now just starting to read everything you just wrote and forgive me but I suck at math and I am not that fast at learning it easily can you please first explain what the question is asking me for part 1 and then explain how to do it?